RFBG Re-Commissioning with AGW2: 18 October 2010 UT

Observer: JHill (LBTO)
Telescope Operator: SAllanson, GBechetti (LBTO)
Telescope Support: MMidkiff (LBTO)
Instrument Support: JMorris (LBTO)

Summary

Lost to weather..........clouds.

We didn't have an Adaptive Secondary Operator, so things would have been problematic if the weather was clear.

Despite clouds, we got lots of closed-dome tasks done related to RFBG wakeup and other developments. We also spent several hours on telescope operator training.

Contacted Runa and Armando in the Italian morning. But we are still covered by clouds.

Plan

In the case of actual clearness we would have been working on the plan: Com201006RFBGPlan
The two main points of emphasis would be:
  • turn on everything at RFBG and make sure it is communicating and operating correctly
  • pursue the astigmatism convergence issue in light of the lenslet rotations seen on MODS AGW. Doug's IDL code has been modified to correct for this rotation, although GCS has not yet been modified.

Preparations

Butch didn't find anything wrong with chillers. See Issue 2935. (more below)

Dan installed new DSP software for the rotators to have self-generated slewing polynomials.

Details

Testing Rotators

We made some closed dome presets to test the rotator code. Dan fixed one problem for us on RFBG, but there are still issues with LFBG for Dan to fix tomorrow. No, we forgot to switch the PCS from right to left. All the rotators seem to be OK.

Testing LBTtools.Observe

We tested the parts of DaveT's new IRAF package LBTtools.Observe that could be tested with AO5. Dave fixed one minor bug for us in takepic and takeseq.

Crashed the IRS server

During the above tests, the IRS server crashed two times with segmentation fault. We are not able to repeat the crashes with the same commands. Called Paul, but he didn't know the answer. See Issue 2937.

Testing Cooling for Mirror Ventilation

We used the suggestion of Butch to turn on both coolant circulation pumps PH0401 and PH0402 manually. This got Mirror Ventilation running OK under similar conditions to last night. See Issue 2935. Later we found that PH0401 is producing 15% less pressure than PH0402.

Debugging Steve's pointnear.pl scripts

We worked for a little while on Steve's Perl script to find a nearby pointing star. We were not able to figure out why the script cannot spawn an IRAF session with a file of commands.

Telescope Operator Training

We continued the activity of simulating errors in the telescope-enclosure systems. Then we send Geno out to debug and fix them. Restarting the new SX air compressor after an E-stop turned out to be more challenging than expected. The "automatic" LED has to be lit before the ECSPLC can turn the compressor on in remote mode.

IRTC Windows PC Security Updates

We installed 40 pending updates on the IRTC Windows PC, rebooted the machine, and verified that IRTC still operated correctly afterwards.

The End

11:05 Sky has cleared, but humidity is 92% and rising. This brings our chances of observing to an official end.

-- JohnHill - 18 Oct 2010
Topic revision: r3 - 18 Oct 2010, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback